# How Aligned are Human Chart Takeaways and LLM Predictions? A Case Study on Bar Charts with Varying Layouts

### v1

@@ 4.2.2 Context affects LLM comparison types more than data
Figure 8 shows four distributions of comparison types induced by the four charts for each layout. Except for the stacked layout, the distributions visually appear more similar across datasets (same shape, different colors) than across contexts (same color, different shapes). Table 4.2.2 confirms that the average TVD between distributions from different contexts is larger than that between distributions from different datasets with the exception of the stacked layout. Therefore, we conclude that variations in LLM comparison types within the same layout is more attributable to context than to data in most cases.
@@ 5 Limitations and Future Work
Task-Specific Alignment: Existing work has demonstrated that user takeaways from visualizations can be task-dependent [20, 24]. In the present study, we focused on eliciting overall takeaways from charts. Future work can additionally test whether the misalignment between LLM outputs and human takeaways also takes place when the visualization is provided in the context of a more specific task, such as low-level analytic tasks [2]. Specifically, because humans tend to be prone to cognitive biases even when completing objective analytic tasks with visualizations, such as estimating correlations [41, 17], it would be interesting to explore the extent to which LLMs are capable of predicting these biases in human behaviors.
Prompting and Finetuning: In Experiment 1, we tested for an optimal configuration of chart specifications, temperature, prompting strategies, and LLM type using the adjacent bar layout. Future work can test for optimal configurations across a wider range of charts, layouts, and datasets with an eye towards better generalizability.
Additionally, the two prompting strategies we tested serve as a starting point to demonstrate the power of prompting. Considering the black-box nature of LLMs, there could be an infinite number of ways to prompt LLMs to obtain even stronger results. Future work could explore the design space of prompting LLMs for human-aligned chart takeaways. We suspect adding elements that highlight aspects of the visualization an LLM should pay special attention to, such as pointing out a design feature that has been manipulated, would increase LLM-human takeaway alignment. Future exploration of the prompt space can also consider manipulating the number of examples provided to the LLMs. While the zero-shot and one-shot settings tested in this work are fast and practical, increasing the number of in-context examples has the potential to better calibrate chart takeaways generated by LLMs in light of the benefits of scaling in-context learning in other domains [1]. Beyond prompting, another exciting opportunity for takeaway alignment is finetuning. By finetuning LLMs on pairs of charts and human-written takeaways, we might be able to improve their ability to generate human-aligned takeaways.
@@ 6 Conclusion
Through a case study on bar charts with different layouts, we provide an initial attempt at understanding the extent to which LLMs are perceptually aware when generating visualization takeaways. In Experiment 1, we identified the optimal configurations to generate chart takeaways. We found that even state-of-the-art LLMs can sometimes struggle to generate semantically diverse and factually accurate takeaways. In Experiment 2, we prompted LLMs to generate human-aligned takeaways across multiple spatial layouts. We found that LLMs are generally not sensitive to bar chart layout like humans are. Nonetheless, providing an example chart and some human takeaways helps align model generations in some cases. In Experiment 3, we observed that, across different data and contexts, while humans generally reported similar comparison types across charts with the same layout, LLMs showed tremendous variability in what they compared. This shows that LLMs are too reliant on context and data when generating takeaways. We further demonstrated that context affects LLM comparison types more than data for most layouts.

### v2

@@ 4.2.2 Context affects LLM comparison types more than data
Figure 8 shows four distributions of comparison types induced by the four charts for each layout. With the exception of the stacked layout, we see that the distributions visually appear more similar across datasets (same shape, different colors) than across contexts (same color, different shapes). Table 4.2.2 confirms that the average TVD between distributions from different contexts are larger than that between distributions from different datasets with the exception of the stacked layout. Therefore, we conclude that LLM comparison type within the same layout is more attributable to context than to data in most cases.
@@ 5 Limitations and Future Work
Task-Specific Alignment: Existing work has demonstrated that user takeaways with visualizations can be task-dependent [20, 24]. In the present study, we focused on eliciting overall takeaways from charts. Future work can additionally test whether the misalignment between LLM outputs and human takeaways also takes place when the visualization is provided in the context of a more specific task, such as low-level analytic tasks [2]. Specifically, because humans tend to be prone to cognitive biases even when completing objective analytic tasks with visualizations, such as estimating correlations [41, 17], it would be interesting to explore the extent to which LLMs are capable of predicting these biases in human behaviors.
Prompting and Finetuning: In Experiment 1, we tested for an optimal configuration of chart specifications, temperature, prompting strategies, and input representation using the adjacent bar chart. Future work can test for optimal configurations across a wider range of charts, layouts, and datasets with an eye towards better generalizability.
Additionally, the two prompting strategies we tested serve as a starting point to demonstrate the power of prompting. Considering the black-box nature of LLMs, there could be an infinite number of ways to prompt LLMs to obtain even stronger results. Future work could explore the design space of prompting LLMs for human-aligned chart takeaways. We suspect adding elements that highlight aspects of the visualization an LLM should pay special attention to, such as pointing a design feature that has been manipulated, would increase LLM-human takeaway alignment. Future exploration of the prompt space can also consider manipulating the number of examples provided to the LLMs. While the zero-shot and one-shot settings tested in this work are fast and practical, increasing the number of in-context examples has the potential to better calibrate chart takeaways generated by LLMs in light of the benefits of scaling in-context learning in other domains [1]. Beyond prompting, another exciting opportunity for takeaway alignment is finetuning. By finetuning LLMs on pairs of charts and human-written takeaways, we might be able to improve their ability to generate human-aligned takeaways.
@@ 6 Conclusion
Through a case study on bar charts with different layouts, we provide an initial attempt to understand the extent to which LLMs are perceptually aware when generating visualization takeaways. In Experiment 1, we identified the optimal configurations to generate chart takeaways. We found that even state-of-the-art LLMs can sometimes struggle to generate semantically diverse and factually accurate takeaways. In Experiment 2, we prompted LLMs to generate human-aligned takeaways across multiple spatial layouts. We found that LLMs are generally not sensitive to bar chart layout like humans are. Nonetheless, providing an example chart and some human takeaways helps align model generations in some cases. In Experiment 3, we observed that, across different data and contexts, while humans generally reported similar comparison types across charts with the same layout, LLMs showed tremendous variability in what they compared. This shows that LLMs are too reliant on context and data when generating takeaways. We further demonstrated that context affects LLM comparison types more than data for most layouts.